From: Jet Propulsion Laboratory
Posted: Friday, September 4, 2009
The most recent spacecraft telemetry was acquired on Sept. 1 from the Deep Space Network tracking complex at Goldstone, California. The Cassini spacecraft is in an excellent state of health and all subsystems are operating normally. Information on the present position and speed of the Cassini spacecraft may be found on the "Present Position" page at: http://saturn.jpl.nasa.gov/operations/present-position.cfm .
Wednesday, Aug. 26 (DOY 238):
A two-day Titan Science Workshop held at the California Institute of Technology began today. Topics on the agenda included presentations and discussions of Titan's gravity, hypsometry, spin rate, temporal changes, surface composition, surface temperature, sea-shoreline elevations, lake distribution, composition and shape, and future plans for observations in the proposed Cassini Extended Extended mission.
This week Cassini science observations included Visual and Infrared Mapping Spectrometer (VIMS) high phase and low phase mosaics of the rings and an occultation of alpOri by the F - D rings. The Ultraviolet Imaging Spectrograph (UVIS) performed an additional low incidence angle occultation. Imaging Science (ISS) imaged Janus in the ring shadow, observed a Telesto close flyby at ~24,700 km and 86 deg phase at closest approach, performed a Titan cloud monitoring activity, looked for shadows of vertical features in the rings, performed a 9.5h radial scan of the main rings, and created a lit face, high phase movie in search of ring spokes. The Composite Infrared Spectrometer (CIRS) also created a low-phase thermal map of the rings and performed a 6h equinox baseline observation. Finally both the Cassini Plasma Spectrometer (CAPS) and Magnetospheric and Plasma Science (MAPS) suite of instruments each performed a 2-hour survey.
Thursday, Aug. 27 (DOY 239):
Non-targeted flybys of Pallene and Telesto occurred today.
In the Significant Events Report for Tuesday, Aug. 18, it was reported that an anomaly occurred in the Spacecraft Office (SCO) Integrated Test Lab (ITL) between a CDS hardware interface and the support equipment. Today SCO reported that the ITL is back up and functioning in dual string mode with the CDS flight spare and engineering model computers. The problem was isolated to a support equipment interface board. The functionally equivalent board was removed from CDS Test Bed 1 and moved to ITL Test Bed 3. SCO is now working on having the problem board refurbished.
Friday, Aug. 28 (DOY 240):
At 6:30 PM, Pacific Daylight Time, JPL employees were notified that due to brush fires in the area, the primary Oak Grove site would be closed to all but essential personnel over the weekend as a precautionary measure. In this case for Cassini, essential personnel included individuals needed for the successful execution of Orbit Trim Maneuver (OTM)-215 on Saturday morning.
Saturday, Aug. 29 (DOY 241):
Due to ongoing issues with the LCROSS spacecraft, it was requested that Cassini give up the beginning of two DSN tracks on DOY 241 and 243. For DOY 241, DSN schedulers were able to obtain a short DSS-55 track to maintain the full OTM-215 uplink window. On DOY-243, the decision was made - with the concurrence of project management - to accept the science data loss. This will result in approximately 45 MB of data that will not be seen on the ground due to the loss of DSS-65 track time.
Orbit Trim Maneuver (OTM) #215 was performed today. This is the cleanup maneuver from the Titan 61 encounter on Aug. 24. The main engine burn began at 7:45 AM PDT. Telemetry immediately after the maneuver showed the burn duration was 2.97 seconds, giving a delta-V of 0.50 m/s. All subsystems reported nominal performance after the maneuver.
Monday, Aug. 31 (DOY 243):
JPL Update: Sunday, August 30, 2009, 4:45 pm Fire conditions around JPL have continued to improve throughout the day, and the Station Fire no longer threatens the facility. However, there has been heavy smoke in the area. To ensure acceptable air quality for the safety of employees, the Laboratory was closed on Monday except for mission-critical personnel. A decision regarding reporting to work on Tuesday will be made by Monday afternoon.
Tuesday, Sept. 1 (DOY 244):
JPL resumed normal operations at 0600 hours Tuesday, September 1.
Science Planning handed over S55 sequence development to Uplink Operations today at the Science and Sequence Update Process (SSUP) kickoff meeting. SSUP takes approximately 10 weeks and is the final sequence development process just prior to uplink and execution of the background sequence and all related products. The first two weeks of DSN allocation beginning on Nov. 13 have been negotiated. S55 DSN allocations should be finalized by Sept. 9.
An image titled "Shadows of Saturn at Equinox" is Astronomy Picture of the Day today. To view the image and caption link to: http://antwrp.gsfc.nasa.gov/apod/ap090901.html
An AACS Periodic Engineering Maintenance (PEM) occurred today. Performed every 90 days, this activity exercises the main engine gimbal actuators, the backup Reaction Wheel Assembly wheel, and scrubs some portions of memory. In the RWA exercise, the wheel is commanded to +100 rpm, -100 rpm, 0 rpm, then turned off.
The Science Forum for S57 was held today. Topics included an overview of science planned for this sequence followed by highlights/unique/highest priority observations provided by the Target Working Teams and Orbiter Science Team leads, with comments from the Investigation Scientists and other instrument team representatives.
Visit the JPL Cassini home page for more information about the Cassini Project: http://saturn.jpl.nasa.gov/
// end //